OpenStudy (anonymous):

3x to third y plus 6x squared y to the third

OpenStudy (anonymous):

HELP ASAP

OpenStudy (solomonzelman):

\[2x^3+6x^2y^3\]this?

OpenStudy (anonymous):

factor x squared plus 6x+9 please???

OpenStudy (anonymous):

u can forget the first one

OpenStudy (solomonzelman):

\[x^2+6x+9~~~~~~~~~~~~~~this?\]

OpenStudy (solomonzelman):

\[(x+3)(x+3)~~~~~~~~~~~~~~~or~~~~~~~~~~~~~~~~~(x+3)^2\]

OpenStudy (anonymous):

yes! and how u found it

OpenStudy (anonymous):

howds u find it

OpenStudy (solomonzelman):

i know that 9 can be written as\[9 \times 1~~~~~~~~~~~or~~~~~~~~~~~3 \times 3\]and 3+3=6 to me it's just a common sense.

OpenStudy (anonymous):

how does adding a constant affect the graph of a quadratic equation

OpenStudy (solomonzelman):

\[(x+3)(x+3)~~~~~~~~~~~~~~~or~~~~~~~~~~~~~~~~~(x+3)^2\]

OpenStudy (anonymous):

how does adding a constant affect the graph of a quadratic equation??

OpenStudy (anonymous):

What is the discontinuity and zero of the function f(x) = the quantity of 2 x squared plus 5 x minus 12, all over x plus 4?